The increase in the volume of 19cm³ of mercury when the temperature rises by 100°¢ is 0.182cm³.what is the cubic expansivity of mercury

1
Expert's answer
2022-03-01T15:18:09-0500

By definition, cubic expansivity is given as follows:


"\alpha = \dfrac{1}{V}\dfrac{\Delta V}{\Delta T} = \dfrac{1}{19cm^3}\cdot \dfrac{0.182cm^3}{100\degree C} \approx 9.6\times 10^{-5}\degree C^{-1}"

Answer. "9.6\times 10^{-5}\degree C^{-1}".